# Interpreting artificial neural networks to detect genome-wide association signals for complex traits

### v1

@@ Abstract
Investigating the genetic architecture of complex diseases is challenging due to the highly polygenic and interactive landscape of genetic and environmental factors. Although genome-wide association studies (GWAS) have identified thousands of variants for multiple complex phenotypes, conventional statistical approaches can be limited by simplified assumptions such as linearity and lack of epistasis models. In this work, we trained artificial neural networks for predicting complex traits using both simulated and real genotype/phenotype datasets. We extracted feature importance scores via different post hoc interpretability methods to identify potentially associated loci (PAL) for the target phenotype. Simulations we performed with various parameters demonstrated that associated loci can be detected with good precision using strict selection criteria, but downstream analyses are required for fine-mapping the exact variants due to linkage disequilibrium, similarly to conventional GWAS. By applying our approach to the schizophrenia cohort in the Estonian Biobank, we were able to detect multiple PAL related to this highly polygenic and heritable disorder. We also performed enrichment analyses with PAL in genic regions, which predominantly identified terms associated with brain morphology. With further improvements in model optimization and confidence measures, artificial neural networks can enhance the identification of genomic loci associated with complex diseases, providing a more comprehensive approach for GWAS and serving as initial screening tools for subsequent functional studies.
@@
Deep learning, interpretability, genome-wide association studies, complex diseases
@@ Introduction
Understanding the genetic component of complex diseases is challenging due to high polygenicity and confounding factors defining these traits. Genome-wide association studies (GWAS) have been crucial for the discovery of thousands of disease-related genomic loci in the past two decades [1], and they continue to be relevant thanks to ever-expanding biobanks with deep phenotype/genotype datasets, allowing the detection of the most minuscule but statistically significant signals. Despite the immense success and broad applications over all types of traits, conventional GWAS methodology still relies on linear models (i.e., logistic and linear regression) for detecting such signals. More recently, linear mixed models have been utilised to incorporate random and fixed effects, especially to account for relatedness and population structure without data pruning [2, 3]. Although such improvements seem to improve GWAS outcomes generally, they do not diverge from the fundamental assumptions of linear and additive genetic models for complex traits. A natural alternative for the nonlinear modelling of such traits would be machine learning, and especially deep learning, since sufficiently large neural networks can approximate any bounded continuous function as defined in the universal approximation theorem [4]. Hence, not surprisingly, various models have been developed for the prediction of disease status from genetic data following the success of deep learning applications in many domains [5, 6]. Despite the state-of-the-art performance of these models in prediction, a main drawback of utilising such complex architectures is the inherent difficulty of interpretability due to the high number of parameters at various abstraction levels. In this context, the development and integration of interpretability methodology are not only important for understanding model biases and improving architectures but also for detecting novel features (e.g., genomic loci, haplotypes, or variants) associated with the target phenotype. Various approaches have been proposed to tackle interpretability for the potential discovery of causal genomic loci, ranging from intrinsically interpretable architectures based on domain knowledge, such as functional gene annotations [7, 8, 9], to post hoc analysis of trained models [10, 11, 12, 13]. Despite these important advancements, multiple challenges remain for the widespread use of interpretable deep learning in genomics. For instance, many of the proposed methods rely on either specific architectures, specific interpretability methods, or a priori biological knowledge. In this work, (i) we propose a general framework for obtaining potentially associated loci (PAL) from neural network models trained for phenotype prediction, (ii) compare the performance of different post hoc and model-agnostic interpretability methods and logistic regression using simulations and (iii) apply our approach to the Estonian Biobank (EstBB) [14] schizophrenia (SCZ) cohort demonstrating its utility for GWAS and reporting potentially novel loci not reported in literature.
@@
We assembled the SCZ cohort from the EstBB dataset so that all patients (ICD codes F20-F29) would have at least one antipsychotic prescription and have their first diagnosis after 2006 with the diagnosis age between 14 and 60. For SCZ cases, we first assembled a set of individuals without any mental, behavioural and neurodevelopmental disorders (no ICD F* diagnosis) and without any antipsychotic prescription. From this filtered set, controls were selected to match the age, sex and BMI distribution of the case cohort with four controls per case using MatchIt [15]. We used exact matching (method = “exact”) for sex and nearest neighbour matching (method = “nearest”) for BMI and age, resulting in 1814 cases (642 male, 1172 female) and 7325 controls (2586 male, 4739 female). In addition, the whole dataset was pruned to eliminate relatedness based on identity by descent (pi_hat $<$ 0.2). We curated the genotype data of the EstBB SCZ case/control cohort derived from the GSA array (Illumina’s GSA-MD-24v1, GSA-MD-24v2, ESTchip1_GSA-MD-24v2, ESTchip2_GSA-MD-24v3, 2022-09-14 snapshot) with a final set of 290,522 SNPs over 22 autosomal chromosomes after removing SNPs unique to at least one version of the array, indels and filtering for bi-allelic regions.
@@
We obtained the simulated phenotypes from the same combined SCZ case/control genotype dataset. Instead of simulating genotypes, we simulated phenotypes to keep genome distribution as realistic as possible. More in detail, let $X$ represent the genotype matrix where $X_{ij}$ is the genotype at position $j$ for individual $i$ with possible values -1 (AA), 0 (AB), 1 (BB); $\beta$ the vector of genetic weights sampled from standard normal distribution $N$(0,1); $\epsilon_{i}$ the noise term sampled from $N$(0, $k\sigma^{2}$) where $k$ is a scaling factor and $\sigma^{2}$ is the variance of the total genetic effects. Then the continuous phenotype $Y$ for individual $i$ is defined as:
@@
In practice, we chose $\tau$ to have approximately the same number of cases ($\sim$1814) as the real SCZ cohort. We simulated a total of 6 scenarios with different noise scaling factors ($k$ = 1, $k$ = 2, $k$ = 3) and two different causal SNP sizes (n = 100 and n = 1000) where all four genetic effect types (dominant, recessive, two-way and three-way interactive) had positions with ratios 5:5:8:2, respectively (i.e., for n = 100: 40, 10, 25, 25 and for n = 1000: 400, 100, 250, 250). Genotypes were represented by -1, 0, 1 instead of 0, 1, 2 to provide equal importance to either allele in a given position (considering the neural network weights are initialised with Kaiming uniform distribution, including negative and positive values uniformly). Nevertheless, our preliminary trials with 0, 1, 2 representation provided similar results in terms of model training.
@@
Models trained for binary trait classification were all feed forward fully connected neural networks with an input layer, two hidden layers and an output layer. The specific architecture consists of an initial dropout layer to randomly mask a proportion of features (p = 0.99) at each batch training, a linear layer with input size = 290,522 and output size = 290, a dropout layer (p = 0.6) followed by ReLU activation, a linear layer with input size = 290 and output size = 29, a dropout layer (p = 0.6) followed by GELU activation, a linear layer with input size = 29 and output size = 1 followed by sigmoid activation function. Training was performed using cross-entropy loss function and Adam optimizer with learning rate = 1e-5 and weight decay = 1e-3 for regularisation. Initially, we tried a weighted loss function to compensate for the class imbalance in the real dataset but this did not work (i.e., no training run where the validation loss decreases), so we instead duplicated the case samples and reduced the sample size of the controls by half with random subsampling to match the case/control numbers [16]. We additionally added random noise $N$(0,0.1) both to class labels (case label = 1, control label = 0) and input genotypes to increase generalizability. All models were trained for 1000 epochs with batch size = 256. All models were coded with python-3.9 and pythorch-1.11 [17].
@@
There is a plethora of different interpretability methods for deep learning models, but we focused on three post hoc and model agnostic approaches. Two of them, gradients w.r.t input (saliency maps, SM) [18] and integrated gradients (IG) [19], are based on gradients and applicable to any differentiable model. SM creation is a commonly used and straightforward approach where gradients w.r.t input are obtained to assess which features in a given input are important for the output probability. For a binary classification neural network model $f$ and input vector $x$, the saliency (with absolute values) for the $j^{th}$ feature is defined as:
@@
As the authors of the method emphasise, the choice of baseline vector is important so that a meaningful interpolation can be achieved in the given domain. In our application, we tested various baselines and decided on using a vector of zeros (where the input is a case genotype) which in preliminary analyses, provided better outcomes compared to (i) [input = case and control genotypes] - [baseline = vector of average allele dosages for the whole dataset] and (ii) [input = case genotypes] - [baseline = control genotypes] settings. In this regard, our choice of genotype representations with -1, 0, 1 might also have been helpful since a vector of zeros would correspond to a completely heterozygous genome, without preference over any allele in a given position. We used Captum-0.6.0 python library [20] for calculating SM and IG.
@@
Here, we opted to measure the rate of change in predictions for the feature importance instead of a performance metric such as accuracy to increase sensitivity.
Using these three methods, we obtained absolute feature importance scores for a set of case samples (both in real and simulated scenarios) for all trained models, averaged the scores over samples and applied min-max normalisation to map the values to [0,1] range. We call this averaged value mean attribution score (MAS). Moreover, the obtained case set consisted of genotypes with neural network predicted probabilities above the 70th percentile. This thresholding was performed mainly for computational ease but also for utilising case samples with prediction values closer to 1. Different percentile thresholds and using all true case genotypes provided similar results in preliminary analyses.
@@
To account for the stochasticity of neural networks and reduce false positive rate, we devised a streamlined method for obtaining potentially associated loci (PAL) from the MAS computed from multiple models trained with different seeds. More specifically, let $A$ be the MAS matrix sized $m\times L$ where $m$ is the number of models (i.e., models trained with different seeds) and $L$ is the length of the genotype. We obtain the average MAS vector $\mu$:
@@
Then we set a significance threshold $\theta$ and obtain weights $w$ based on the number of occurrences above the $\theta$ threshold among $m$ models trained with different seeds:
@@
In practice, we chose $m=10$ for models trained with different seeds to be able to perform a large number of trials with realistic computational resources. Since there were multiple PAL in high LD detected by models trained with different seeds, we also considered these positions ($r^{2}>0.5$) between different models as common signals between models to increase detection power both for $PAL_{Common}$ and $PAL_{AMAS}$. Furthermore, we set $\theta$ to 99.99 percentile (strict) and 99.95 percentile (relaxed) thresholds. Although the choice of $\theta$ is essentially arbitrary, the strict threshold produces 30 PAL and the relaxed threshold 145 PAL, which are plausible numbers for keeping the false positive rate low considering the expected high polygenicity of schizophrenia [22]. An overview of our proposed approach is provided in Fig 1.
@@
Logistic regression
To compare our method to conventional GWAS, we performed logistic regression (LR) using all case (n = 1814) and control (n = 7325) samples (both in real and simulated scenarios) using the top 3 principal components (PCs) from principal component analysis (PCA) as covariates. We obtained the significance threshold by dividing p-value = 0.05 by the number of tests (i.e., number of SNPs: 290,522). We used python-3.9 and statsmodels 0.14.1 [23] for the code. Furthermore, we performed an association analysis using REGENIE v2.2.4 [3] on real data with the same case/control genotypes via the methodology described by Pujol Gualdo et al. [24], except for the filtering steps to retain exactly the same individuals and positions as LR.

### v2

@@ Abstract
Investigating the genetic architecture of complex diseases is challenging due to the multifactorial and interactive landscape of genomic and environmental influences. Although genome-wide association studies (GWAS) have identified thousands of variants for multiple complex traits, conventional statistical approaches can be limited by simplified assumptions such as linearity and lack of epistasis in models. In this work, we trained artificial neural networks to predict complex traits using both simulated and real genotype-phenotype datasets. We extracted feature importance scores via different post hoc interpretability methods to identify potentially associated loci (PAL) for the target phenotype and devised an approach for obtaining p-values for the detected PAL. Simulations with various parameters demonstrated that associated loci can be detected with good precision using strict selection criteria. By applying our approach to the schizophrenia cohort in the Estonian Biobank, we detected multiple loci associated with this highly polygenic and heritable disorder. There was significant concordance between PAL and loci previously associated with schizophrenia and bipolar disorder, with enrichment analyses of genes within the identified PAL predominantly highlighting terms related to brain morphology and function. With advancements in model optimization and uncertainty quantification, artificial neural networks have the potential to enhance the identification of genomic loci associated with complex diseases, offering a more comprehensive approach for GWAS and serving as initial screening tools for subsequent functional studies.
@@
Deep learning, interpretability, genome-wide association studies, complex traits, schizophrenia
@@ Introduction
Understanding the genetic component of complex diseases is challenging due to high polygenicity and influence of confounding factors. Genome-wide association studies (GWAS) have been crucial for the discovery of thousands of disease-related genomic loci in the past two decades [1], and they continue to be relevant due to ever-expanding biobanks with extensive phenotype and genotype data, which enable the detection of the most minuscule but statistically significant signals. Despite the immense success and broad applications across various complex traits, conventional GWAS methodology still relies on linear models (i.e., logistic and linear regression) for detecting signals. More recently, linear mixed models have been utilised to incorporate random and fixed effects, primarily to account for relatedness and population structure without data pruning [2, 3]. While these advancements have generally improved GWAS outcomes, they do not diverge from the core assumptions of linearity and additivity in genetic models for complex traits. A natural alternative for nonlinear modelling of such traits would be machine learning, and particularly deep learning, since sufficiently large neural networks can approximate any bounded continuous function as defined in the universal approximation theorem [4]. Hence, not surprisingly, various models have been developed for the prediction of disease status from genetic data following the success of deep learning applications in many domains [5, 6]. Despite the state-of-the-art performance in prediction, a main drawback of employing such complex architectures is the inherent difficulty of interpretability due to the high number of parameters at various abstraction levels. In this context, the development and integration of interpretability methodology is not only important for understanding model biases and improving architectures, but also for detecting novel features (e.g., genomic loci, haplotypes, and specific genetic variants) associated with the target phenotype. Various approaches have been proposed to tackle interpretability for the discovery of trait-associated loci, ranging from intrinsically interpretable architectures based on domain knowledge, such as functional gene annotations [7, 8, 9], to post hoc analysis of trained models [10, 11, 12, 13, 14]. Despite these important advancements, multiple challenges remain for the widespread use of interpretable deep learning in genomics. For instance, many of the proposed methods rely on either specific architectures, specific interpretability methods, or a priori biological knowledge. In this work, we (i) propose a general framework for identifying potentially associated loci (PAL) from neural network models trained for phenotype prediction using genome-wide SNP array data, (ii) introduce an approach for obtaining p-values for the detected PAL, (iii) compare the performance of various post hoc and model-agnostic interpretability methods using simulations, and (iv) apply our approach to the schizophrenia (SCZ) cohort at Estonian Biobank (EstBB) [15, 16] demonstrating its utility for GWAS with potentially novel loci not reported in literature.
@@
We assembled the SCZ cohort from the EstBB dataset so that all patients (ICD codes F20-F29) would have at least one antipsychotic prescription and have their first diagnosis after 2006 with the diagnosis age between 14 and 60 [17]. As controls for the SCZ cases, we first assembled a set of individuals without any mental, behavioural and neurodevelopmental disorders (no ICD F* diagnosis) and no record of antipsychotic prescriptions. From this filtered set, controls were selected to match the age, sex and BMI distribution of the case cohort with four controls per case using MatchIt [18]. We used exact matching (method = “exact”) for sex and nearest neighbour matching (method = “nearest”) for BMI and age, resulting in 1814 cases (642 male, 1172 female) and 7325 controls (2586 male, 4739 female). In addition, the whole dataset was pruned to eliminate relatedness based on identity by descent (pi_hat $<$ 0.2) [19, 20]. We curated the genotype data of the EstBB SCZ case/control cohort derived from the GSA array (Illumina’s GSA-MD-24v1, GSA-MD-24v2, ESTchip1_GSA-MD-24v2, ESTchip2_GSA-MD-24v3, 2022-09-14 snapshot). Genotype data quality control was performed according to best practices. Specifically, individuals with call rate $<95\%$, who deviated ±±±±3SD from the samples’ heterozygosity rate mean or showed mismatch between heterozygosity of the X chromosome and sex based on phenotype data were excluded; all AT and GC SNPs, invariable SNPs, SNPs showing potential traces of batch bias, poor cluster separation results and inconsistent allele frequency among any of the EstBB genotyping experiments were removed. The final set contained 290,522 SNPs across 22 autosomal chromosomes after further removing SNPs unique to at least one version of the array, indels and filtering for bi-allelic regions.
@@
We obtained simulated phenotypes from the same combined SCZ case/control genotype dataset. Instead of simulating genotypes, we simulated phenotypes to keep genome distribution as realistic as possible. More in detail, let $X$ represent the genotype matrix where $X_{ij}$ is the genotype at position $j$ for individual $i$ with possible values -1 (AA), 0 (AB), 1 (BB); $\beta$ the vector of genetic weights sampled from standard normal distribution $N$(0,1); $\epsilon_{i}$ the noise term sampled from $N$(0, $k\sigma^{2}$) where $k$ is a scaling factor and $\sigma^{2}$ is the variance of the total genetic effects. Then the continuous phenotype $Y$ for individual $i$ is defined as:
@@
In practice, we chose $\tau$ to have approximately the same number of cases ($\sim$1814) as the real SCZ cohort. We simulated a total of 6 scenarios with different noise scaling factors ($k$ = 1, $k$ = 2, $k$ = 3) and two different numbers of causal SNPs (n = 100 and n = 1000) where all four genetic effect types (dominant, recessive, two-way and three-way interactive) had positions with ratios 5:5:4:6, respectively (i.e., for n = 100: 25, 25, 20, 30 and for n = 1000: 250, 250, 200, 300), yielding equal numbers of causal SNPs for dominant/recessive versus interactive effect types. Genotypes were represented by -1, 0, 1 instead of 0, 1, 2 to provide equal importance to either allele in a given position (considering that the neural network weights are initialised with Kaiming uniform distribution, including negative and positive values uniformly). Nevertheless, our preliminary trials with 0, 1, 2 representation provided similar results in terms of model training.
@@
Models trained for binary trait classification were all feedforward fully connected neural networks with an input layer, two hidden layers and an output layer. The specific architecture consists of an initial dropout layer to randomly mask a proportion of features (p = 0.99) at each batch training; a linear layer with input size = 290,522 and output size = 290; a dropout layer (p = 0.6) followed by ReLU activation; a linear layer with input size = 290 and output size = 29; a dropout layer (p = 0.6) followed by GELU activation; a linear layer with input size = 29 and output size = 1 followed by sigmoid activation function. This equates to total 84,260,139 trainable parameters. Training was performed using cross-entropy loss function and Adam optimizer with learning rate = 1e-5 and weight decay = 1e-3 for regularisation. Initially, we implemented a weighted loss function to address the class imbalance in the dataset; however, this approach was ineffective (i.e., none of the training runs showed a reduction in validation loss). As an alternative, we duplicated the case samples and reduced the control sample size by half through random subsampling to balance the case/control ratio [21]. We additionally added random noise $N$(0,0.1) both to class labels (case label = 1, control label = 0) and input genotypes to increase generalizability. All models were trained for 1000 epochs with batch size = 256 using 150 case and 150 control samples as validation. All models were coded with python-3.9 and pythorch-1.11 [22].
@@
While there is a plethora of different interpretability methods for deep learning models, we focused on three post hoc and model-agnostic approaches to be able to produce a general framework suitable for most architectures. Two of them, gradients w.r.t input (saliency maps, SM) [23] and integrated gradients (IG) [24], are based on gradients and applicable to any differentiable model. SM creation is a commonly used and straightforward approach where gradients w.r.t input are obtained to assess which features in a given input are important for the output probability. For a binary classification neural network model $f$ and input vector $x$, the saliency (with absolute values) for the $j^{th}$ feature is defined as:
@@
As the authors of the method emphasise, the choice of baseline vector is important so that a meaningful interpolation can be achieved in the given domain. In our application, we tested various baselines and decided on using a vector of zeros (where the input is a case genotype) which in preliminary analyses, provided better outcomes (i.e., lower FP and higher TP counts) compared to (i) [input = case and control genotypes] - [baseline = vector of average allele dosages for the whole dataset] and (ii) [input = case genotypes] - [baseline = control genotypes] settings. In this regard, our choice of genotype representations with -1, 0, 1 might also have been helpful, since a vector of zeros would correspond to a completely heterozygous genome, without preference over any allele in a given position. We used Captum-0.6.0 python library [25] for calculating SM and IG.
@@
Here, we opted to measure the rate of change in predictions for the feature importance instead of a performance metric such as accuracy, as this approach increases sensitivity. This is because each feature is expected to have only a small contribution without causing substantial alteration in the output prediction.
Using these three methods, we obtained absolute feature importance scores for the case samples (both in real and simulated scenarios) for all trained models, applied L1 normalization across genomic positions and averaged the scores over samples. We call this averaged value mean attribution score (MAS). L1 normalization was performed to transform the score of each feature for a given sample into a measure of contribution, enabling comparisons between samples.
@@
To account for the stochasticity of neural networks and reduce false positive rate, we devised a streamlined method for obtaining potentially associated loci (PAL) from MAS computed from multiple models trained with different seeds. More specifically, let $A$ be the MAS matrix sized $m\times L$ where $m$ is the number of models (i.e., models trained with different seeds) and $L$ is the length of the genotype. We obtain the average MAS vector $\mu$:
@@
Then we set a threshold $\theta$ and obtain weights $w$ based on the number of occurrences above the $\theta$ threshold among $m$ models trained with different seeds:
@@
In practice, we chose $m=10$ for models trained with different seeds to be able to perform a large number of tests with realistic computational resources. Since multiple PAL in high LD were detected by models trained with different seeds, we considered these positions ($r^{2}>0.5$) across models as common signals to increase detection power. Furthermore, we set the threshold $\theta$ to 99.99 percentile (strict) and 99.95 percentile (relaxed). Although the choice of $\theta$ is essentially arbitrary, the strict threshold yields 30 PAL and the relaxed threshold 145 PAL, which are plausible numbers for maintaining a low false positive rate, given the high polygenicity of schizophrenia [27]. An overview of our proposed approach is provided in Fig 1.
Obtaining statistical significance
The main approach with the $\theta$ threshold weighting is designed for mitigating random effects of stochasticity (i.e., a few “rogue” models presenting high scoring non-associated loci) but it does not provide a statistical framework for controlling false positives. A common approach for this purpose would be a complete permutation-based method but since this would be computationally infeasible in our case, we instead devised a method to obtain p-values based on the general assumption that higher attribution values are more likely to indicate a true signal. Initially, we trained 10 models with permuted labels (i.e., shuffled binary disease status) and obtained null MAS for each, using the same steps as described previously. Since MAS is calculated as the absolute mean over multiple samples, a reasonable expectation (due to central limit theorem) for the null distribution would be a folded normal distribution with $\mu=0$, which was approximately satisfied in IG null MAS at the tail region of interest (S1 Fig). We estimated the scale parameter $\sigma=4.2\mathrm{e}{-6}$ from the pooled null MAS and sampled 10 different MAS vectors from the fitted half-normal distribution. To compute AMAS for both simulated and real datasets, we reordered each of the 10 sampled null MAS vectors to match the exact ranking of each of the observed 10 MAS vectors. This procedure emulates dependencies both within (i.e., due to LD and neural network connectivity) and between vectors (i.e., due to shared training data and architecture). We then averaged sampled null MAS vectors and calculated AMAS as previously described. This sampling procedure was repeated 100 times, and p-values were derived by counting the number of sampled AMAS values exceeding the target AMAS. The pseudocode for the method is provided below.
Algorithm 1 Deriving $p$-values for AMAS
0:  MAS matrix $A$ of size $m\times L$ (with $m$ models and $L$ loci), significance threshold $\theta$, and number of sampled iterations $R$.
1:  Step 1: Compute Observed AMAS
2:  Compute the average MAS vector: $\mu_{j}=\frac{1}{m}\sum_{a=1}^{m}A_{aj}$ for $j=1,\dots,L$.
3:  Compute weights: $w_{j}=\frac{1}{m}\sum_{a=1}^{m}\mathbb{I}(A_{aj}>\theta)$, where $\mathbb{I}$ is the indicator function.
4:  Compute observed AMAS: $\mathrm{AMAS}_{j}=w_{j}\cdot\mu_{j}$ for all $j$ such that $\mu_{j}>\theta$.
5:  Step 2: Generate Null Distribution
6:  for $r=1$ to $R$ do
7:     Sample $m$ null MAS vectors from $\text{foldedNormal}(0,\sigma)$, forming the null MAS matrix $A_{\text{null}}^{(r)}$ of size $m\times L$.
8:     for $a=1$ to $m$ do
9:        Sort the null MAS values for model $a$ to match the rank order of the observed MAS values for the same model:
$A_{a,:,\text{null}}^{(r)}\leftarrow\text{sort}(A_{a,:,\text{null}}^{(r)},\text{ by order of }A_{a,:})$
10:     end for
11:     Compute the average null MAS vector: $\mu_{j}^{(r)}=\frac{1}{m}\sum_{a=1}^{m}A_{aj,\text{null}}^{(r)}$.
12:     Compute null weights: $w_{j}^{(r)}=\frac{1}{m}\sum_{a=1}^{m}\mathbb{I}(A_{aj,\text{null}}^{(r)}>\theta)$.
13:     Compute null AMAS: $\mathrm{AMAS}_{j}^{(r)}=w_{j}^{(r)}\cdot\mu_{j}^{(r)}$ for all $j$ such that $\mu_{j}^{(r)}>\theta$.
14:  end for
15:  Step 3: Compute $p$-values
16:  Combine all null AMAS values across all loci and iterations into a single pooled null distribution.
17:  for $j=1$ to $L$ do
18:     Compute $p$-values as:
$p_{j}=\frac{\sum_{r=1}^{R}\sum_{k=1}^{L}\mathbb{I}(\mathrm{AMAS}_{k}^{(r)}>\mathrm{AMAS}_{j})}{R\cdot L}$
19:  end for
20:  return  $p_{j}$ for all $j$.
@@
Conventional GWAS
To compare our method to conventional linear GWAS, we initially performed logistic regression (LR) using all case (n = 1814) and control (n = 7325) samples (both in real and simulated scenarios) using the top 3 principal components (PCs) from principal component analysis (PCA) as covariates. LR without PC covariates did not affect the results (as there is no or limited population structure in the EstBB dataset) (S2 Fig). We used python-3.9 and statsmodels 0.14.1 [28] for the LR code. Furthermore, we also performed an association analysis using a state-of-the-art LMM-based tool REGENIE v2.2.4 [3] on real data via the exact protocol described by Pujol Gualdo et al. [29], except for the filtering steps to retain the same individuals and genomic positions as LR.
@@
Functional analyses
For PAL detected with the SCZ dataset, we defined genes in the ±$\pm 100$kb region of the detected SNPs as positionally mapped genes. For expression analyses and eQTL mapping, we used the variant-gene association dataset (GTEx_Analysis_v8_eQTL) of the open access Adult Genotype Tissue Expression (GTEx) Project [30]. For all SNPs in PAL, we assessed the significance of expression levels (significance defined as pval_nominal $<$ pval_nominal_threshold in significant variant-gene pairs dataset from GTEx *.signif_variant_gene_pairs.txt.gz files) in all brain tissues, and reported eQTL mapped genes for each PAL as genes for which the expression was altered significantly. To assess the functionality of positionally mapped genes in the detected PAL, we used the GENE2FUNC tool in FUMA [31], which leverages tissue-specific expression levels and phenotype enrichment based on the GWAS catalog [32]. Additionally, to assess the function of each variant, we used the g:Profiler g:SNPense tool [33].
Since we did not perform fine-mapping and multiple SNPs in detected PAL can be due to the same signal, we presented findings only related to our lead SNPs in the literature. More specifically, (i) we extracted all our positionally mapped genes from GWAS Catalog, (ii) identified the SNPs mapped to these genes in the catalog and (iii) presented LD between SNPs reported to be associated with a neuropsychiatric disorder and our lead SNPs. This is not an exhaustive approach and rather conservative, since GWAS Catalog often does not include all significant SNPs in a detected region, and SNPs we detected could potentially be related to other SNPs, especially considering our non-imputed array framework.